How many grams of glocose are needed to lower the vapor pressre of water by 100.0 torr?

Other question:

A solution is prepared by mixing 23.5mL of carbon tetrachloride and 35.8 mL of 2,5-dichlorohexane. What is the expected vapor pressure of this solution? What is the moele fraction of carbon tetrachloride in the vapor above the solution? The density of carbon tetrachloride is 1.594 g/mL and its vapor pressure is 91 mmHg. The density of 2,5-dichlorohexane is 1.0g/mL and its vapor pressure is 2.3 mmHg.
1
Expert's answer
2018-05-24T07:37:28-0400
At 54C vapor pressure of water is 112.5 torr.
P(solution) = x(solvent)*P(solvent)
12.5 = x*112.5
x = 0.11
n(C6H12O6) = 0.88 mol
m(C6H12O6) = 0.88mol*180g/mol = 158.4 g
m(CCl4) = 23.5mL*1.594g/mL = 37.459g
m(C6H12Cl2) = 35.8g
n(CCl4)= 37.459g/154g/mol = 0.24mol
x(CCl4)=0.24/0.24+0.23 = 0.51
P(solution) = 0.51*91 = 46.41mmHg